# Distilling Large Language Models for Network Active Queue Management

### v1

@@ Abstract
The growing complexity of network traffic and demand for ultra-low latency communication require smarter packet traffic management. Existing Deep Learning-based queuing approaches struggle with dynamic network scenarios and demand high engineering effort. We propose AQM-LLM, distilling Large Language Models (LLMs) with few-shot learning, contextual understanding, and pattern recognition to improve Active Queue Management (AQM) [RFC 9330] with minimal manual effort. We consider a specific case where AQM is Low Latency, Low Loss, and Scalable Throughput (L4S) and our design of AQM-LLM builds on speculative decoding and reinforcement-based distilling of LLM by tackling congestion prevention in the L4S architecture using Explicit Congestion Notification (ECN) [RFC 9331] and periodic packet dropping. We develop a new open-source experimental platform by executing L4S-AQM on FreeBSD-14, providing interoperable modules to support LLM integration and facilitate IETF recognition through wider testing. Our extensive evaluations show L4S-LLM enhances queue management, prevents congestion, reduces latency, and boosts network performance, showcasing LLMs’ adaptability and efficiency in uplifting AQM systems.
@@ I Introduction
Since their inception in the 1990s, Active Queue Management (AQM) algorithms have relied on mathematical models to develop optimal rule-based control systems. Early algorithms, such as Random Early Detection (RED) [1], used queue length as the primary metric to track and manage network buffers. Modern AQMs, like Controlled Delay (CoDel) [2], Proportional Integral Controller Enhanced (PIE) [3], and Common Applications Kept Enhanced (CAKE), have shifted to queue delay as the key metric, providing more direct control over user-perceived latency and improving network performance. See Low Latency, Low Loss, and Scalable Throughput (L4S) [RFC 330] [^1] https://www.rfc-editor.org/rfc/rfc9330.html [4] and Bell Labs recent findings for details [5].
Despite these advancements, rule-based AQMs face significant challenges in today’s dynamic and heterogeneous networking environments, which include advances in WiFi, 5G/6G, optical fibers, and satellites. Emerging network environments experience rapid fluctuations in link quality, traffic patterns, and interference levels, often outpacing the adaptive capabilities of formulaic approaches. Addressing such dynamics and heterogeneous network challenges using rule-based methods requires the development of complex, computationally intensive rules that demand periodic (re)engineering and continuous parameter tuning based on specific network dynamics [^2] https://www.bell-labs.com/institute/white-papers/l4s-low-latency-low-loss-and-scalable-throughput/, accessed, 25 Jan 2025. [6].
Deep learning offers a promising alternative, shifting the focus toward learning-based AQMs that minimize reliance on predefined mathematical models [7]. Supervised learning (SL) methods are frequently employed for prediction tasks, such as traffic classification and bandwidth estimation, while reinforcement learning (RL) is utilized for decision-making problems, including congestion control, adaptive bitrate streaming, and cloud cluster job scheduling [8]. In AQM, deep neural networks (DNNs) and RL have been applied to packet traffic prediction and optimization tasks [5]. Compared to rule-based systems, learning-based methods (e.g., DNN, RL, etc.) excel at adapting dynamically to evolving network conditions and leveraging data-driven insights to handle complex scenarios effectively [5]. However, challenges persist, including scalability, robustness, and computational efficiency, which must be addressed to fully realize the potential of ML-driven AQMs, which are summarized as follows.
1. Designing and tuning ML-based AQM models demands significant effort, including architecture design, hyperparameter optimization, and algorithm selection, requiring extensive experimentation and computational resources, which can hinder deployment in dynamic networks.
2. ML-based AQM often struggles to generalize across varying network conditions, with models trained on static traffic performing poorly under real-world fluctuations, limiting their reliability compared to traditional rule-based systems.
3. Continuous retraining of ML-based AQM is essential to adapt to evolving networks, requiring diverse, high-quality datasets and extensive preprocessing, which increases operational costs and complicates commercial deployment.
@@ I-A LLMs over AQM: Challenges & Distillation
Leveraging Large Language Models (LLMs) like GPT, Opt, and Llama [9, 10] into AQM can address the high costs of developing task-specific DNNs for solving networking challenges [11]. Pretrained on vast text datasets, LLMs excel in contextual understanding, reasoning, and generalization [10] making them ideal for tasks like AQM. Their attention mechanism enables dynamic analysis of traffic parameters and dependencies, allowing proactive congestion management and adaptability to evolving network conditions. The distillation of LLM for AQM offers a universal framework to improve the Internet with minimal modifications, which can perhaps be achieved by adopting new RL ideas for the distillation of knowledge [10, 12]. However, key challenges include:
i) Diverse Inputs: AQM data includes time series (e.g., queue delay) and scalar values (e.g., burst allowance), requiring specialized encoders.
ii) Inference Latency: Token-by-token output generation causes delays, disrupting AQM’s rapid update cycles (45 ms). Hallucinated outputs further exacerbate the delays.
iii) Distillation Costs: Distilling LLMs for AQM requires time-consuming fine-tuning, especially in decision-making tasks using reinforcement learning, and demands significant computational resources for large models.
While adopting LLMs can unify and enhance AQM, the above mentioned challenges must be addressed for effective AQM-LLM design and deployment.
I-B Our Design and Contributions
We propose AQM-LLM and develop it over L4S. L4S-LLM is an adapted LLM framework developed for efficient congestion prevention with L4S. L4S-LLM leverages a foundational LLM model for decision-making in the L4S queues with minimal manual modifications, by incorporating three key modules:
State Encoder: A multifaceted encoder processes diverse network data (e.g., queue delay, packet drops) into token-like embeddings, enabling the LLM to understand and utilize network metrics effectively.
L4S-LM Head: Replacing the default language model head, the L4S-LM head maps LLM outputs directly to congestion prevention actions (enqueue, drop, or mark) in a single inference, reducing response time and preventing hallucinations.
Data-Driven Low-Rank L4S Adaptation (proposed LoRA): This module finetunes the LLM using RL with distillation over the pre-collected data from existing algorithms, eliminating lengthy real-time interactions. By employing Low-Rank Adaptation (LoRA), finetuning costs are drastically reduced, with a 64% decrease in GPU memory usage and a 15.1% reduction in training time.
@@ I-B Our Design and Contributions
    Development of L4S-LLM with a multi-faceted encoder, specialized L4S-LM head, and the proposed LoRA scheme for improved RL-based distillation and finetuning of the L4S-LLM.
@@
L4S-LM Head Design & Development. Conventional LM heads rely on iterative multiround inference to generate valid responses, resulting in substantial latency. To overcome this limitation, we propose the novel L4S-LM head, which generates a valid response in a single inference round by predicting a probability distribution over possible actions. This design streamlines the response generation process, significantly accelerating inference. Drawing inspiration from speculative decoding, the L4S-LM head advances beyond sequential token-by-token generation by simultaneously evaluating multiple potential outcomes, enhancing both computational efficiency and scalability (without degrading any accuracy). Furthermore, L4S-LM offers a robust framework to improve AQM performance while reducing manual effort and resource consumption.
@@ I-C Literature
While these methods improve adaptability, they face challenges like high engineering effort, computational cost, and resource-intensive training. Their reliance on continuous learning limits generalization across diverse network environments, hindering practical deployment in real-world settings. LLMs [19] like ChatGPT, PaLM, Llama2, and OPT are advanced deep neural networks built on the Transformer architecture, with widespread applications across diverse domains. LLMs process input and output as token sequences, converting text into embeddings and predicting subsequent tokens using an auto-regressive mechanism. The self-attention mechanism allows LLMs to focus on relevant parts of the input, enabling them to handle complex language patterns and nuances effectively.
Trained on vast text datasets, LLMs demonstrate exceptional capabilities in dialogue, reasoning, and code generation. Fine-tuning these models for domain-specific tasks enhances their few-shot learning abilities, allowing them to generalize quickly with minimal examples. This adaptability and knowledge transfer highlight their potential for solving complex tasks efficiently, making them invaluable for optimizing AQM schemes. Inspired by these advancements, our proof-of-concept adapts LLMs for AQM network management within the L4S architecture. We develop novel ways of applying LLMs to the DualPI2 AQM algorithm in FreeBSD, leveraging their reasoning and few-shot learning abilities to generalize across dynamic, unseen conditions.
@@ II Research Design
DualPI2 uses ECN (RFC 3168 [20]) to mark packets during congestion, resorting to packet drops only under severe congestion. ECN leverages two bits in the IP header to signal congestion using four codepoints, with CE (1) indicating congestion. Upon receiving CE-marked packets, the receiver informs the sender via an ECE flag, prompting the sender to adjust congestion control and respond with a CWR flag to acknowledge the notification. This mechanism mitigates delays, reduces packet loss, and minimizes retransmission, leading to lower latency, improved throughput, and decreased head-of-line blocking [21].
@@
In this paper, we develop an approach that uses the extraordinary capabilities of LLM to prevent congestion in the L4S DualPI2 by deciding the appropriate action to be taken in that specific scenario. Mainly, its job is to decide between packet dropping, packet enqueueing, and most importantly, ECN notification. LLMs are trained on billions of parameters that absorb extensive knowledge, allowing them to show extraordinary abilities such as planning, pattern mining, problem-solving, and generalization to unseen conditions.
Our objective is to utilize the LLMs’ prediction capabilities to enhance the efficacy and foresight in identifying network congestion within the AQM framework. If there is congestion ahead, the LLM will inform TCP using the ECN notification or will directly drop the packets if the delay is excessive. This will enable TCP congestion control algorithms to take preemptive action. By accurately predicting congestion events within L4S architecture, TCP Prague is able to mitigate any unnecessary delays, retransmission, and packet loss during network congestion.
Figure 1: L4S-LLM Architecture
Figure 2: L4S-LLM Coupling: Interaction between different components of L4S and LLM.
II-A Architecture of AQM-LLM
Figure 3: L4S-LLM Encoder: Details of the components
In this section, we detail the design of AQM-LLM, a framework that leverages Large Language Models to optimize AQM tasks, including packet enqueueing, dropping, and ECN marking. AQM-LLM comprises three core components: (i) the L4S-LLM State Encoder, (ii) the L4S-LLM Head, and (iii) Data-driven Low-Rank L4S Adaptation. Below, we provide an overview of these components, focusing on the L4S-LLM State Encoder.
The L4S-LLM State Encoder processes task-specific information such as queue size, delay, and other AQM metrics. It converts raw input data from various modalities into a format compatible with LLMs by projecting these inputs into a token space tailored to the specific model. This ensures seamless integration with LLMs, where the token space size adapts to the underlying model.
As illustrated in Figure 3, the encoder consists of two main components: (a) Feature Encoder: Extracts meaningful features from raw AQM metrics such as queue delay and size; (b) Linear Projector: Maps these features into token-like embeddings aligned with the LLM’s input requirements. Together, these components enable efficient processing of diverse AQM inputs, ensuring compatibility and effectiveness in LLM-driven queue management.

### v3

@@ Abstract
We propose AQM-LLM, a framework that distills Large Language Models for Active Queue Management in modern networks. Current Deep Learning-based queuing approaches struggle with dynamic network scenarios and require extensive engineering effort. Our approach leverages LLMs’ contextual understanding and pattern recognition capabilities to enhance Low Latency, Low Loss, and Scalable Throughput (L4S) architecture with minimal manual intervention. The L4S-LLM implementation incorporates three key innovations: (1) an improved state encoder that processes diverse network metrics into token embeddings, (2) an extended LLM head that generates single-inference congestion decisions, and (3) a data-driven Low-Rank Adaptation scheme that reduces trainable parameters considerably while maintaining very high accuracy. Our open-source FreeBSD-14 implementation demonstrates significant improvements in queue management metrics, with reduced delay variability and enhanced bandwidth utilization across both DCTCP and UDP Prague protocols. Comprehensive evaluations show L4S-LLM effectively balances computational efficiency with network performance, establishing a foundation and novel direction for intelligent, adaptive queuing in next-generation networks.
@@ I Introduction
Since the 1990s, Active Queue Management (AQM) algorithms have employed mathematical models for optimal rule-based control systems. Early algorithms like Random Early Detection (RED)[1] utilized queue length as the primary buffer management metric. Modern AQMs including Controlled Delay (CoDel)[2], Proportional Integral Controller Enhanced (PIE)[3], and Common Applications Kept Enhanced (CAKE) have transitioned to queue delay-based metrics. This shift enables direct latency control and enhanced network performance. The Low Latency, Low Loss, and Scalable Throughput (L4S) framework [RFC 9330][4] exemplifies these advances [^1] https://www.rfc-editor.org/rfc/rfc9330.html. Recent Bell Labs findings validate these approaches [5].
Rule-based AQMs encounter critical limitations in contemporary heterogeneous networks. WiFi, 5G/6G, optical fiber, and satellite environments exhibit rapid link quality fluctuations. Traffic patterns and interference levels change faster than formulaic adaptation capabilities. Dynamic heterogeneous challenges necessitate complex, computationally intensive rules. Continuous parameter tuning and periodic re-engineering, perhaps by employing machine learning (ML), become mandatory [^2] https://www.bell-labs.com/white-papers/l4s/, accessed 15 Aug 2025[6].
Deep learning provides a paradigm shift toward learning-based AQMs. These systems minimize dependence on predefined mathematical models [7]. Supervised learning handles traffic classification and bandwidth estimation tasks. Reinforcement learning addresses congestion control, adaptive bitrate streaming, and cloud scheduling [8]. Deep neural networks (DNNs) and RL optimize packet traffic prediction in AQM contexts [5]. Learning-based methods demonstrate superior dynamic adaptation compared to rule-based systems. Data-driven insights enable effective handling of complex scenarios [5]. Nevertheless, critical challenges impede ML-driven AQM deployment:
(1) Design Complexity: ML-based AQM model development demands substantial effort. Architecture design, hyperparameter optimization, and algorithm selection require extensive experimentation. Computational resource requirements hinder dynamic network deployment.
(2) Generalization Limitations: ML-based AQMs struggle with cross-condition generalization. Models trained on static traffic exhibit poor real-world performance. Reliability remains inferior to traditional rule-based systems.
(3) Retraining Overhead: Continuous ML-based AQM retraining is essential for network evolution adaptation. Diverse, high-quality datasets and extensive preprocessing increase operational costs. Commercial deployment complexity escalates significantly.
@@ I-A LLMs over AQM: Challenges & Distillation
Leveraging Large Language Models (LLMs) like GPT, Opt, and Llama [9, 10] into AQM can address the high costs of developing task-specific DNNs for solving networking challenges [11, 12]. Pretrained on vast text datasets, LLMs excel in contextual understanding, reasoning, and generalization [10] making them ideal for tasks like AQM. Their attention mechanism enables dynamic analysis of traffic parameters and dependencies, allowing proactive congestion management and adaptability to evolving network conditions. The distillation of LLM for AQM offers a universal framework to improve the Internet with minimal modifications, which can perhaps be achieved by adopting new RL ideas for the distillation of knowledge [10, 13]. However, key challenges include:
(1) Diverse Inputs: AQM data includes time series (e.g., queue delay) and scalar values (e.g., burst allowance), requiring specialized encoders.
(2) Inference Latency: Token-by-token output generation causes delays, disrupting AQM’s rapid update cycles (45 ms). Hallucinated outputs further exacerbate the delays.
(3) Distillation Costs: Distilling LLMs for AQM requires time-consuming fine-tuning, especially in decision-making tasks using reinforcement learning, and demands significant computational resources for large models. While adopting LLMs can unify and enhance AQM, the above mentioned advances must be developed for effective AQM-LLM design and deployment.
I-B Our Design Details of L4S-LLM and Contributions
We propose AQM-LLM and develop and implement L4S-LLM as a proof of concept. L4S-LLM is an adapted LLM framework developed for efficient congestion prevention on the top of L4S. L4S-LLM leverages a foundational LLM model for decision-making in the L4S queues with minimal manual modifications, by incorporating three key modules:
(1) State Encoder: A multifaceted encoder processes diverse network data (e.g., queue delay, packet drops) into token-like embeddings, enabling the LLM to understand and utilize network metrics effectively.
(2) L4S-LLM Head: Replacing the default language model head, the L4S-LLM head maps LLM outputs directly to congestion prevention actions (enqueue, drop, or mark) in a single inference, reducing response time and preventing hallucinations.
(3) Data-Driven Low-Rank L4S Adaptation (proposed LoRA): This module finetunes the LLM using RL with distillation over the pre-collected data from existing algorithms, eliminating lengthy real-time interactions. By employing Low-Rank Adaptation (LoRA), finetuning costs are drastically reduced, with a 64% decrease in GPU memory usage and a 15.1% reduction in training time.
@@ I-B Our Design Details of L4S-LLM and Contributions
    Development of L4S-LLM with a multi-faceted encoder, specialized L4S-LLM head, and the proposed LoRA scheme for improved RL-based distillation and finetuning of the L4S-LLM.
@@
L4S-LLM Head Design & Development. Conventional LLM heads rely on iterative multiround inference to generate valid responses, resulting in substantial latency. To overcome this limitation, we propose the novel L4S-LLM head, which generates a valid response in a single inference round by predicting a probability distribution over possible actions. This design streamlines the response generation process, significantly accelerating inference. Drawing inspiration from speculative decoding, the L4S-LLM head advances beyond sequential token-by-token generation by simultaneously evaluating multiple potential outcomes, enhancing both computational efficiency and scalability (without degrading any accuracy). Furthermore, L4S-LLM offers a robust framework to improve AQM performance while reducing manual effort and resource consumption.
Figure 1 demonstrates the integration of LLM intelligence with L4S network stack. Observe in Figure 1 how LLM components analyze real-time network telemetry and traffic patterns to dynamically optimize L4S parameters, including dual-queue AQM management, ECN marking thresholds, and congestion control algorithms. A bidirectional feedback loop enables LLM to process network performance data while simultaneously providing intelligent parameter adjustments and predictive congestion management to the existing L4S stack, resulting in enhanced network efficiency and reduced latency through LLM-driven L4S.
Figure 1: L4S-LLM Coupling: Interaction between different components of L4S and LLM, illustrating how the model integrates with the existing network stack.
Temporal Modeling Framework. Our temporal modeling framework employs a dual-stage architecture to capture queue dynamics across multiple timescales. The initial stage utilizes 1D convolutional neural networks (Algorithm 1) with varying kernel sizes to extract multi-scale temporal features from sequential network metrics, providing efficient feature extraction before transformer processing.
The transformer component employs self-attention across an empirically optimized context window of network activity. Ablation studies revealed this configuration maximizes performance while maintaining generalizability. Shorter context windows exhibited performance degradation, while extended windows showed no significant improvement and increased computational complexity. To mitigate temporal overfitting, we implemented three regularization techniques:
  1. 1.
    Protocol-diverse training corpus: Incorporating heterogeneous network traces from multiple congestion control algorithms
  2. 2.
    Stochastic temporal augmentation: Applying random offsets and noise to temporal sequences
  3. 3.
    Temporal dropout: Implementing random masking of timestep information during training
Our findings highlight the complementary strengths of learning-based and classical approaches. While L4S-LLM offers superior performance in most scenarios, classical algorithms provide valuable safety guarantees and interpretability that inform our ongoing work to combine these strengths through formal verification and explicit incorporation of classical control principles. Empirical validation demonstrates the efficacy of this approach, with the model maintaining high accuracy on previously unseen network conditions. The data-driven reinforcement learning pipeline optimizes for discounted cumulative returns rather than immediate rewards, improving long-horizon performance compared to myopic optimization.
@@ I-C Literature
While these methods improve adaptability, they face challenges like high engineering effort, computational cost, and resource-intensive training. Their reliance on continuous learning limits generalization across diverse network environments, hindering practical deployment in real-world settings. LLMs [20] like ChatGPT, PaLM, Llama2, and OPT are advanced deep neural networks built on the Transformer architecture, with widespread applications across diverse domains. LLMs process input and output as token sequences, converting text into embeddings and predicting subsequent tokens using an auto-regressive mechanism. The self-attention mechanism allows LLMs to focus on relevant parts of the input, enabling them to handle complex language patterns and nuances effectively. Nevertheless, the LLM Driven networking has potential in several networking aspects including network function virtualization and large scale simulations [21, 22, 23].
There are some notable approaches for adapting LLMs to networking domains [24, 25]. Supervised Fine-Tuning (SFT) [26] enables efficient domain adaptation with minimal data while preserving instruction-following capabilities, but exhibits safety degradation that increases with dataset size and is particularly sensitive to structured formats like lists and tables [27]. Continual Pre-Training (CPT) [28] provides deeper domain knowledge integration and superior performance on complex tasks, but suffers from severe safety degradation when not followed by safety-focused instruction tuning [28, 29]. To address these safety concerns, several methods have been proposed that maintain safety guardrails while preserving task performance. These include SafeInstruct [30], which interleaves safety-aligned QA pairs during fine-tuning; SafeLoRA [31], which selectively adapts only layers exhibiting harmful behavior; and SafeMERGE [32], which merges fine-tuned layers with those of known safe models.
@@ II Research Design
DualPI2 uses ECN (RFC 3168 [33]) to mark packets during congestion, resorting to packet drops only under severe congestion. ECN leverages two bits in the IP header’s DiffServ field to signal congestion with four states. Table I shows the details. Non-ECT (00) indicates endpoints don’t support ECN, so routers drop packets during congestion. ECT(1) (01) and ECT(0) (10) show endpoints are ECN-capable, allowing routers to mark packets instead of dropping them. CE (11) indicates the packet experienced congestion and is set by routers to notify the receiving endpoint. [^5] In our FreeBSD implementation, we have configured ECN through the net.inet.tcp.ecn.enable sysctl parameter. The default setting activates ECN solely for incoming connections requesting it, but can be modified to apply to all connections or turned off completely.
Upon receiving CE-marked packets, the receiver informs the sender via an ECE flag, prompting the sender to adjust congestion control and respond with a CWR flag to acknowledge the notification. This mechanism mitigates delays, reduces packet loss, and minimizes retransmission, leading to lower latency, improved throughput, and decreased head-of-line blocking [34]. An abstract view of our L4S-LLM Architecture with complete pipeline from raw network metrics to AQM decisions, are shown in Figure 2 with the data flow through the state encoder, transformer model, and L4S-LLM head. We discuss the components of the L4S-LLM Architecture in details as follows.
@@
II-A Kernel Log Collection and Format
Our approach begins with collecting raw network metrics directly from the FreeBSD kernel’s DualPI2 implementation. These logs capture the complete state of the AQM system at each decision point. Figure 3 shows an example of the raw kernel log in the format (observe the values in the highlighted dashed rectangles): queue type, qdelay reference, tupdate, max burst, max ecn threshold, alpha coefficient, beta coefficient, flags, burst allowance, drop probability, current queue delay, previous queue delay, accumulated probability, measurement start time, average dequeue time, dequeue count, status flags, total packets, total bytes, queue length, length in bytes, total drops, packet length, dequeue action.
Figure 2: L4S-LLM Architecture: The complete pipeline from raw network metrics to AQM decisions, showing the data flow through the state encoder, transformer model, and L4S-LLM head.
Figure 3: Snapshot of Raw Kernel Log
Figure 4: Snapshot of our processed experience pool
Each log entry represents a single packet processing event with timestamps and key metrics. The final field in the dashed rectangle, action, indicates the decision made by the original DualPI2 algorithm: 0 for enqueue, 1 for drop, and 2 for ECN marking. See Table I for details. These logs are collected using a custom kernel module seamlessly that hooks into the DualPI2 decision path without affecting normal operation.
II-B Experience Pool Generation
The raw kernel logs are processed into structured experience pools for training and evaluating our L4S-LLM model. Each entry contains a state vector of 8 network condition features, an action decision (0: enqueue, 1: drop, 2: ECN mark), a computed reward (balancing throughput and latency), and a Boolean flag refer to as Done indicating episode termination (0 except at transmission end). Figure 4 shows the snapshot of the processed experience pool. The experience pools are stored in a structured JSON format to facilitate efficient loading during model training.
We demonstrate the L4s-LLM pipeline overview from data collection to LLM head in Table II. Table III displays all metric data obtained from the kernel along with their descriptions. Custom logging functions were implemented within the kernel to accurately capture data for analysis, offline data cleaning, and preparation for generating the LLM experience pool.
TABLE II: Pipeline Stages and Overview
Stage Description Example
Kernel Log Collection Real-time AQM metrics collected from kernel-level code during packet processing Queue Type = 1, Queue Delay = 50ms, Drop Prob = 0.02
Experience Pool Generation Kernel logs transformed into training/evaluation pools with state, action, reward, and done flags State: $f_{1},f_{2},\ldots,f_{n}]$, Action: $\{0,1,2\}$, Reward: throughput, Done: $\{0,1\}$
Tokenization Network data converted into discrete tokens for LLM processing Network state $\rightarrow$ [queue_delay, drop_prob, pkt_len, …]
Embedding Tokens mapped to high-dimensional vectors matching LLM input space queue_delay $\rightarrow[0.2,0.7,0.1,\ldots]$, action $\rightarrow[0.3,0.1,-0.1,\ldots]$
RL Adaptation L4S-LLM fine-tuned using rewards to optimize AQM decisions $s_{t}=(qdelay_{t},pdrop_{t},lenb_{t})$, $R_{t}=\frac{pkt\_len_{t}}{qdelay_{t+1}}$
LoRA Parameter-efficient training using low-rank matrices while freezing LLM weights $\Delta W=A\cdot B$, where $A,B$ are low-rank matrices
L4S Head Maps LLM outputs to congestion actions (enqueue, drop, mark) $Action=\arg\max(P(enqueue),P(drop),P(mark))$
TABLE III: Collected Kernel-Level Metrics from FreeBSD DualPI2 AQM
Variable Definition
queue_type Queue type identifier (0: Classic, 1: L4S)
burst_allowance Maximum allowed burst size before congestion control activation
drop_probability Probability of dropping a packet during congestion
current_queue_delay Current queue delay experienced by packet
accumulated_probability Accumulated probability for drop decisions
length_in_bytes Current queue length in bytes
packet_length Length of current processed packet
(a) Kernel logs utilized by L4S-LLM
Variable Definition
qdelay_reference Target queue delay reference for AQM algorithms
tupdate Update interval for queue state variables
max_burst Maximum allowable burst duration
max_ecn_threshold Upper ECN marking threshold
alpha_coefficient PI controller proportional gain
beta_coefficient PI controller integral gain
flags Internal AQM status flags
previous_queue_delay Previous measurement cycle queue delay
measurement_start_time Current measurement interval timestamp
average_dequeue_time Mean packet dequeue time
dequeue_count Packets dequeued in current interval
status_flags Operational state indicators
total_packets Total packets processed
total_bytes Total bytes processed
queue_length Current queue length in packets
total_drops Total packet drops in interval
dequeue_action Decision (0:enqueue, 1:drop, 2:ECN mark)
(b) Non-utilized Kernel log parameters
II-C Tokenization and Embedding Process
Unlike traditional NLP applications where text is tokenized into discrete units, our network metrics require specialized processing to be compatible with LLM architectures. We implement a custom tokenization process that converts numerical network metrics into a format suitable for transformer models.
We present a novel approach to network congestion control by adapting transformer-based language models for sequential decision-making in L4S environments. The architecture fundamentally reframes network state transitions as a sequence modeling problem, leveraging the autoregressive capabilities of pre-trained LLMs. Our tokenization process transforms heterogeneous network telemetry data into a unified sequential representation suitable for transformer processing. Rather than traditional linguistic tokens, the system constructs sequences from temporal network observations, where each timestep $t$ generates a multi-modal token comprising: (1) return-to-go $R_{t}$ representing cumulative future rewards, (2) decomposed state features $\mathbf{s}_{t}=[s_{t,1},s_{t,2},\ldots,s_{t,7}]$ capturing network conditions, and (3) action $a_{t}\in\{0,1,2\}$ representing packet handling decisions.
The sequence construction follows the following pattern:
$\mathcal{S}=[R_{t},s_{t,1},s_{t,2},\ldots,a_{t},R_{t+1},s_{t+1,1},\ldots,a_{t+1},\ldots]$ (1)
which creates a causal dependency structure where states predict subsequent actions, aligning with the autoregressive nature of transformer architectures.
Furthermore, the proposed embedding mechanism employs separate linear projection layers for each modality to map heterogeneous input spaces into a unified $d$-dimensional embedding space matching the pre-trained LLM’s hidden size. State features undergo domain-specific transformations through dedicated embedding matrices $\mathbf{W}_{\text{state},i}\in\mathbb{R}^{d\times d_{\text{state}}}$ for $i\in\{1,\ldots,7\}$, while actions and returns are projected via $\mathbf{W}_{\text{action}}\in\mathbb{R}^{d\times 1}$ and $\mathbf{W}_{\text{return}}\in\mathbb{R}^{d\times 1}$ respectively.
Temporal information is incorporated through learned positional embeddings $\mathbf{W}_{\text{time}}\in\mathbb{R}^{(T+1)\times d}$, which are additively combined with all modality embeddings
$\text{embed}(\mathbf{x}_{t})=\mathbf{W}_{\text{modality}}(\mathbf{x}_{t})+\mathbf{W}_{\text{time}}(t)$ (2)
where $\mathbf{x}_{t}$ represents the input at timestep $t$.
The concatenated embedding sequence undergoes layer normalization before being fed into the transformer backbone, bypassing traditional tokenization by directly providing input embeddings rather than discrete tokens. This preserves the continuous nature of network telemetry while leveraging pre-trained attention patterns. The transformer processes sequences through multi-head self-attention mechanisms, capturing complex temporal dependencies and cross-modal interactions between network states, actions, and rewards. The attention mechanism computes weighted representations by calculating similarity scores between query, key, and value matrices, applying softmax normalization, and producing context-aware embeddings. The final hidden states from state embedding positions are extracted and passed through a linear classification head to generate action logits, treating congestion control as a sequence-to-sequence mapping where network history predicts optimal packet handling decisions.
II-D Architecture of AQM-LLM